A coffee machine is regulated so that the amount of drink dispensed is approximately normally distributed with a standard deviation of 15 mm. Find the 95% confidence interval for the mean of all coffee drinks dispensed by this machine if the random sample of 36 drinks had an average content of 225 ml.


1
Expert's answer
2021-05-07T09:59:45-0400

Critical value at 0.05 significance level z = Β±1.96

The 95% confidence interval for the mean is

"225 - (1.96)\frac{15}{\sqrt{36}} < μ < 225 + (1.96) \frac{15}{\sqrt{36}} \\

225 - 4.9 < μ < 225 + 4.9 \\

220.1 < μ < 229.9"